OpenStudy (trojanpoem):

Find the relation used to calculate emf (average) for dynamo in 3/4 period

OpenStudy (trojanpoem):

@Michele_Laino

OpenStudy (michele_laino):

here we have to start from the equation which expresses the f.e.m. E(t) in function of time t

OpenStudy (trojanpoem):

\[emf = -N \frac{ d \phi }{ dt }\]

OpenStudy (michele_laino):

I mean this function: \[E\left( t \right) = {E_0}\sin \left( {\omega t + \phi } \right)\]

OpenStudy (michele_laino):

where: the period T is such that: \[T = \frac{{2\pi }}{\omega }\]

OpenStudy (trojanpoem):

Now ?

OpenStudy (michele_laino):

by definition, the requested average value is given by the subsequent formula: \[\Large \frac{1}{{\frac{{3T}}{4}}}\int_0^{\frac{{3T}}{4}} {{E_0}\sin \left( {\omega t + \phi } \right)\;dt} \]

OpenStudy (trojanpoem):

Now we will get the integration, right ?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

for simplicity we can assume \[\Large \phi = 0\]

OpenStudy (trojanpoem):

Well, nice trick.

OpenStudy (trojanpoem):

\[1/3T/4 \int\limits_{0}^{3T/4} E _{0}\sin(\omega t) dt\] That's what we have.

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now we change variable, namley: \[\omega t = \tau \] where \tau is the new variable

OpenStudy (michele_laino):

so after a simple integration, we get: \[\Large \frac{1}{{\frac{{3T}}{4}}}\int_0^{\frac{{3T}}{4}} {{E_0}\sin \left( {\omega t} \right)\;dt} = - \frac{{4{E_0}}}{{3\omega T}}\left. {\cos \left( {\omega t} \right)} \right|_0^{\frac{{3T}}{4}} = ...?\]

OpenStudy (trojanpoem):

\[\frac{ -4 }{ 3 } E_{0} 1/2\pi \cos(wt) \]

OpenStudy (michele_laino):

no, you have to evaluate the function \cos(\omega*t), between 0 and 3*T/4

OpenStudy (trojanpoem):

Hmm, I am confused.

OpenStudy (michele_laino):

\[\Large \left. {\cos \left( {\omega t} \right)} \right|_0^{\frac{{3T}}{4}} = \cos \left( {\omega \frac{{3T}}{4}} \right) - \cos \left( {\omega \cdot 0} \right) = ...?\]

OpenStudy (michele_laino):

keep in mind that: \[\Large T = \frac{{2\pi }}{\omega }\]

OpenStudy (trojanpoem):

cos(3 * 2* pi / 4 ) - cos 0 = cos(3pi/2) - cos(0) = -1

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now you have to compute this multiplication: \[\Large - \frac{{4{E_0}}}{{3\omega T}} \times \left( { - 1} \right) = ...?\]

OpenStudy (trojanpoem):

Can I replace omega T with 2 pi ?

OpenStudy (michele_laino):

yes!

OpenStudy (trojanpoem):

2/3pi * E_{0}

OpenStudy (michele_laino):

that's right!

OpenStudy (trojanpoem):

so now we will assume that the number of turns of the dynamo is N so it will be 2/3pi NE what else ?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (trojanpoem):

what more ?

OpenStudy (michele_laino):

N are the turns on the rotor of the alternator. we have finished!

OpenStudy (trojanpoem):

so the final relation is \[\frac{ 2 }{ 3 \Pi}NE _{0}\] can we replace E_{0} with BAomega ? and reaplce omega with 2piv so it can be 2/3pi * 2pi * v * B *A * N = 4/3 NBAv ?

OpenStudy (michele_laino):

I got this: \[\Large \frac{{2NBA\omega }}{{3\pi }}\]

OpenStudy (trojanpoem):

And we can replace the omega with 2 pi (nu) , as nu = 1/T

OpenStudy (michele_laino):

ok!